# Novel View Synthesis using DDIM Inversion

### v1

@@ Abstract
Synthesizing novel views from a single input image is a challenging task. It requires extrapolating the 3D structure of a scene while inferring details in occluded regions, and maintaining geometric consistency across viewpoints. Many existing methods must fine‑tune large diffusion backbones using multiple views or train a diffusion model from scratch, which is extremely expensive. Additionally, they suffer from blurry reconstruction and poor generalization. This gap presents the opportunity to explore an explicit lightweight view translation framework that can directly utilize the high-fidelity generative capabilities of a pretrained diffusion model while reconstructing a scene from a novel view. Given the DDIM-inverted latent of a single input image, we employ a camera pose-conditioned translation U-Net, TUNet, to predict the inverted latent corresponding to the desired target view. However, the image sampled using the predicted latent may result in a blurry reconstruction. To this end, we propose a novel fusion strategy that exploits the inherent noise correlation structure observed in DDIM inversion. The proposed fusion strategy helps preserve the texture and fine-grained details. To synthesize the novel view, we use the fused latent as the initial condition for DDIM sampling, leveraging the generative prior of the pretrained diffusion model. Extensive experiments on MVImgNet demonstrate that our method outperforms existing methods. Code is available.
@@
Introduction
Novel view synthesis is a fundamental task in computer vision and graphics. Remarkable works such as NeRFs (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) are extensively used in 3d scene understanding. Several works improve upon these foundational works. However, their dependence on scene-level optimization and the need for a dense set of views limit usability. Diffusion models (Rombach et al. 2022; Podell et al. 2023), have gained significant traction for the task of novel view synthesis (Elata et al. 2025; Tang et al. 2025). A classical approach is to fine-tune these models on 3D datasets along with a module that encodes the 3D geometry into the architecture (Tang et al. 2023; Liu et al. 2024; Long et al. 2024; Huang et al. 2024; Bourigault and Bourigault 2024; Gao et al. 2024). However, the generated outputs lack consistency in multiview reconstruction as the generation is not entirely controllable and results in images of inadequate quality, and often creates blurry results for long-range viewpoint reconstruction.
DDIM (Song, Meng, and Ermon 2020) proposed a deterministic inversion “DDIM Inversion”, which sequentially adds noise to an image to obtain a noisy latent. The noisy latent can be retraced to the original image using DDIM sampling. This latent encapsulates the signal and the noise that contribute to the mean and variance, which changes the distribution of the noise latent at each inversion time. Previous works such as (Garibi et al. 2024; Mokady et al. 2023) try to optimize or configure this noise representation to better align with the given task. (Staniszewski, Kuciński, and Deja 2025) study inversion noise in detail and claim that the DDIM inversion latent space is less manipulative, which makes direct interpolation with this noise latent difficult for tasks such as novel view synthesis and editing.
This paper proposes a method to generate a novel view from a given input image and camera parameters. Our pipeline works entirely in the DDIM-inverted latent space. We first learn to map an input view latent to a target latent using a translation U‑Net called TUNet. This mapping only approximates a coarse-grained version of the target view. This is due to the fact that diffusion models exhibit spectral bias and favor low-frequency components (Choi et al. 2022). In order to induce the high frequency components, we introduce a novel noisy latent fusion strategy. Notably, we use pretrained diffusion model, and only train a lightweight latent‑space translation network, TUNet, for view transformation. We perform extensive experiments in diverse settings, and show that our work extends to unseen categories as well as out of domain images obtained from the web. Sample results are shown in Figure 1. We claim the following key contributions:
@@ Introduction
Related Work
Neural Radiance Field: Neural field approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al. 2021), use learnable functions to map 3D spatial coordinates and viewing directions to volumetric density and color. These models synthesize novel views by performing volumetric rendering via ray marching through the learned scene representation. NeRF has demonstrated that high-quality novel views can be rendered when trained on a dense set of input views.
While recent extensions such as PixelNeRF (Yu et al. 2021), IBRNet (Wang et al. 2021), MultiDiff (Müller et al. 2024), and others (Henzler et al. 2021; Liu et al. 2022; Wu et al. 2023) aim to perform view synthesis from fewer input views, they often suffer in regions with missing or occluded content. Because these models make deterministic predictions without explicit uncertainty modeling, the generated output tends to average over ambiguities, leading to blurry and less plausible reconstruction in unobserved regions.
Gaussian Splatting: 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023; Peng et al. 2024; Zheng et al. 2025) represent scenes using a set of anisotropic 3D Gaussians. Gaussian Splatting methods are deterministic and depend heavily on accurate multi-view geometry or densely sampled camera poses (Li et al. 2024). When applied in sparse-view or single-view settings, they often fail to generate plausible content in unseen regions because they lack generative priors.
In contrast, our work targets novel view synthesis given only a single input image and a target camera pose. This setting spans both short and long-range viewpoint changes. Under such conditions, methods such as NeRF (Mildenhall et al. 2021) and 3DGS (Kerbl et al. 2023) struggle to extrapolate effectively from a single image, even when augmented with generative guidance as in (Szymanowicz et al. 2024; Tewari et al. 2023).
Transformers: NViST (Jang and Agapito 2024) adopts a transformer-based encoder-decoder architecture (Vaswani et al. 2017; Dosovitskiy et al. 2021) to predict a radiance field from a single image, enabling novel view synthesis via NeRF-style volumetric rendering. However, NViST suffers from loss of fine details due to aggressive downsampling (by a factor of 12), and it struggles to synthesize long-range viewpoints (when the target frame is more than 15 frames away from the input frame in a 30‑frame sweep).
@@
Pretrained Diffusion Models: MVDiffusion (Tang et al. 2023), Zero123++ (Shi et al. 2023), SyncDreamer (Liu et al. 2024), Wonder3D (Long et al. 2024), EpiDiff (Huang et al. 2024), BoostDream (Yu et al. 2024), MVDiff (Bourigault and Bourigault 2024), CAT3D (Gao et al. 2024), Magic-Boost (Yang et al. 2024), Cycle3D (Tang et al. 2025), use a pretrained or finetuned diffusion model. MVDiffusion modifies the Stable Diffusion architecture by introducing a cross-branch attention mechanism, known as correspondence-aware attention (CAA), to model inter-view dependencies. SyncDreamer constructs a view frustum feature volume from all the target noisy views and injects these into the pretrained denoising Unet using depth-wise attention layers. EpiDiff (Huang et al. 2024) integrates an attention module guided by epipolar constraints into the intermediate and decoding stages of the U-Net, enabling the model to capture generalized epipolar geometry across views.
@@
The latent text space of diffusion models has also been explored for view synthesis Burgess, Wang, and Yeung-Levy (2024). However, it is typically limited to small-scale datasets and is not directly applicable to scene-level novel view synthesis tasks.
@@
Training Diffusion Model from Scratch: Several recent works train diffusion models from scratch for novel view synthesis, including Tseng et al. (2023), Photometric-NVS (Yu et al. 2023a), DiffDreamer (Cai et al. 2023), GIBR (Anciukevičius et al. 2024), and (Henderson et al. 2024). Photometric-NVS (Yu et al. 2023a) introduces a two-stream latent diffusion architecture that independently processes the source and noisy target views, while exchanging information via pose-conditioned cross-attention mechanisms. GIBR (Anciukevičius et al. 2024) models 3D scenes using IB-planes and trains the diffusion process directly in pixel space, enabling learning of a joint distribution over multi-view observations and camera poses.
@@
Spectral Behavior of Diffusion
In (Kingma and Gao 2023; Falck et al. 2025), authors study the spectral behavior of diffusion. The forward diffusion (Ho, Jain, and Abbeel 2020) process is given by:
@@
where $\mathbf{x}_{0}$ is the clean latent, $\mathbf{x}_{t}$ is the latent corresponding to the timestep $t$ and $\bar{\alpha}_{t}$ is the scaling factor. High-frequency components, representing fine details, are degraded more rapidly and prior to low-frequency components during the forward diffusion process (Kingma and Gao 2023; Falck et al. 2025). This property is also consistent in the reverse process of diffusion. As shown in Choi et al. (2022), diffusion models inherently favor lower frequencies, which implies that more emphasis must be placed on modelling high-frequency details. In addition, the noise component is often observed to deviate from a standard multivariate Gaussian distribution (Staniszewski, Kuciński, and Deja 2025). At later iterations of inversion, the noise encapsulates high-frequency information of the image and is high in variance. The signal variance decreases with inversion time, and the predicted noise’s variance increases with inversion time. The effective DDIM Inversion(Song, Meng, and Ermon 2020) iteration is:
$\mathbf{x}_{t+1}=$ $\underbrace{\left(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}}_{\text{signal / mean},\,\mathbf{z}^{\text{inv}}_{\mu,t+1}}$ (2)
$\hskip 40.00006pt+\underbrace{\sqrt{1-\alpha_{t+1}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}_{\text{noise / variance},\,\mathbf{z}^{\text{inv}}_{\sigma,t+1}},$
where $\mathbf{x}_{t+1}$ is the noisy latent at timestep $t+1$. $\mathbf{x}_{t}$ is the latent at timestep $t$, which is less noisy than $\mathbf{x}_{t+1}$. $\alpha_{t}$ and $\alpha_{t+1}$ are parameters controlling the noise schedule at timesteps $t$ and $t+1$, derived from the noise variance schedule. $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the predicted noise at timestep $t$, estimated by the diffusion U-Net during the reverse process.
@@
The signal/mean in Equation (2) is the coarse-grained image representation on which we perform the view translation. In addition, the noise/variance in Equation (2) encodes image-specific features that are recovered during the denoising process (Staniszewski, Kuciński, and Deja 2025). For the task of novel-view synthesis, this noise/variance can be used to induce high-frequency details into the view-transformed latent, which can then be fed to DDIM sampling. Based on the aforementioned discussion, we formalize two things for the task of novel view synthesis:
@@
DDIM-inverted Latents
Let $\mathbf{z}^{\text{inv}}_{t}$ be the DDIM-inverted latent. If we use this latent at $t=T$, we may see that the DDIM sampled image deviates from the input image, especially when we do it in fewer DDIM steps (Bao et al. 2025; Zeng, Suganuma, and Okatani 2025; Feng et al. 2024). Thus, we fix $t$ = 600 and get our DDIM inverted noisy initial latent in 30 DDIM steps. Further, the signal/mean term $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ of Equation (2) is the diffusion network’s estimate of the clean latent which we obtain at $t$ by denoising $\mathbf{z}^{\text{inv}}_{t}$ according to the diffusion score model $\boldsymbol{\epsilon}_{\theta}$. This signal/mean term is what we feed into TUNet for view translation. We visualise the reconstructed image corresponding to signal/mean term in Figure 3. We observe that the reconstructed image primarily comprises of the low‑frequency components of the input image. Therefore, in order to impose high-frequency information, we make use of the noise/variance term $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ from Equation (2) that re‑injects the predicted noise at the level $t$. We utilize a pretrained latent diffusion model (LDM) (Rombach et al. 2022) as our generative prior, which we use to compute the DDIM-inverted latents. We first describe the TUNet model, followed by the Fusion Strategy. As we fix timestep $t$ at 600, we drop the subscript $t$ while representing signal/mean and noise/variance terms: $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ and $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ in Equation (2) from now on.
Figure 3: Left: Mean of the DDIM inverted latent at tt=600. Latent is decoded using VAE for visualization. Right: Original $\times512$ image.
TUNET Architecture
TUNet is a U-Net  (Ronneberger, Fischer, and Brox 2015) inspired encoder-decoder architecture designed to predict the DDIM-inverted latent’s mean representation of the target view. TUNet introduces cross attention between an input or reference view and a target view, enabling effective feature transfer between viewpoints. The architecture is conditioned on both camera parameters and class embeddings at multiple stages to preserve geometric consistency and semantic integrity.
Input and Conditioning
@@
Encoder (Down Blocks)
@@
Bottleneck and Decoder (Mid + Up Blocks)
@@
Cross-Attention Module
A cross-attention mechanism is integrated in the mid and up blocks, enabling information flow from the reference to the target view using ray information and latent feature alignment. Let $\mathbf{r_{\text{ref}}}$ denote the ray embeddings of the reference view and $\mathbf{r_{\text{tar}}}$ denote the ray embeddings of the target view. We use standard ray parameterization as in NeRF (Mildenhall et al. 2021) to compute ray origins and directions for camera pose encoding to get $\mathbf{r_{\text{ref}}}$ and $\mathbf{r_{\text{tar}}}$. Let $\mathbf{z^{\text{inv}}_{\text{ref},\mu}}$ be the DDIM-inverted latent mean of the reference image, and $\mathbf{f_{\text{tar}}}$ be the intermediate target feature maps at the cross-attention block.

### v2

@@ Abstract
Synthesizing novel views from a single input image is a challenging task. It requires extrapolating the 3D structure of a scene while inferring details in occluded regions, and maintaining geometric consistency across viewpoints. Many existing methods must fine‑tune large diffusion backbones using multiple views or train a diffusion model from scratch, which is extremely expensive. Additionally, they suffer from blurry reconstruction and poor generalization. This gap presents the opportunity to explore an explicit lightweight view translation framework that can directly utilize the high-fidelity generative capabilities of a pretrained diffusion model while reconstructing a scene from a novel view. Given the DDIM-inverted latent of a single input image, we employ a camera pose-conditioned translation U-Net, TUNet, to predict the inverted latent corresponding to the desired target view. However, the image sampled using the predicted latent may result in a blurry reconstruction. To this end, we propose a novel fusion strategy that exploits the inherent noise correlation structure observed in DDIM inversion. The proposed fusion strategy helps preserve the texture and fine-grained details. To synthesize the novel view, we use the fused latent as the initial condition for DDIM sampling, leveraging the generative prior of the pretrained diffusion model. Extensive experiments on MVImgNet and RealEstate10K demonstrate that our method outperforms existing methods. The code is available at https://github.com/Visual-Conception-Group/ddim_nvs .
@@
1 Introduction
Novel view synthesis is a fundamental task in computer vision and graphics. Remarkable works such as NeRFs [30] and 3DGS [23] are extensively used in 3d scene understanding. Several works improve upon these foundational works. However, their dependence on scene-level optimization and the need for a dense set of views limit usability. Diffusion models [35, 33], have gained significant traction for the task of novel view synthesis [10, 44]. A classical approach is to fine-tune these models on 3D datasets along with a module that encodes the 3D geometry into the architecture [43, 27, 29, 21, 3, 14]. However, the generated outputs lack consistency in multiview reconstruction as the generation is not entirely controllable and results in images of inadequate quality, and often creates blurry results for long-range viewpoint reconstruction.
DDIM [40] proposed a deterministic inversion “DDIM Inversion”, which sequentially adds noise to an image to obtain a noisy latent. The noisy latent can be retraced to the original image using DDIM sampling. This latent encapsulates the signal and the noise that contribute to the mean and variance, which changes the distribution of the noise latent at each inversion time. Previous works such as [15, 31] try to optimize or configure this noise representation to better align with the given task. [41] study inversion noise in detail and claim that the DDIM inversion latent space is less manipulative, which makes direct interpolation with this noise latent difficult for tasks such as novel view synthesis and editing.
This paper proposes a method to generate a novel view from a given input image and camera parameters. Our pipeline works entirely in the DDIM-inverted latent space. We first learn to map an input view latent to a target latent using a translation U‑Net called TUNet. This mapping only approximates a coarse-grained version of the target view. This is due to the fact that diffusion models exhibit spectral bias and favor low-frequency components [7]. In order to induce the high frequency components, we introduce a novel noisy latent fusion strategy. Notably, we use pretrained diffusion model, and only train a lightweight latent‑space translation network, TUNet, for view transformation. We perform extensive experiments in diverse settings, and show that our work extends to unseen categories as well as out of domain images obtained from the web. Sample results are shown in Figure 1. We claim the following key contributions:
@@ 1 Introduction
2 Related Work
Neural Radiance Field: Neural field approaches, such as Neural Radiance Fields (NeRF) [30], use learnable functions to map 3D spatial coordinates and viewing directions to volumetric density and color. These models synthesize novel views by performing volumetric rendering via ray marching through the learned scene representation. NeRF has demonstrated that high-quality novel views can be rendered when trained on a dense set of input views.
While recent extensions such as PixelNeRF [59], IBRNet [49], MultiDiff [müller2024multidiffconsistentnovelview], and others [18, 28, 54] aim to perform view synthesis from fewer input views, they often suffer in regions with missing or occluded content. Because these models make deterministic predictions without explicit uncertainty modeling, the generated output tends to average over ambiguities, leading to blurry and less plausible reconstruction in unobserved regions.
Gaussian Splatting: 3D Gaussian Splatting (3DGS) [23, 32, 65] represent scenes using a set of anisotropic 3D Gaussians. Gaussian Splatting methods are deterministic and depend heavily on accurate multi-view geometry or densely sampled camera poses [26]. When applied in sparse-view or single-view settings, they often fail to generate plausible content in unseen regions because they lack generative priors.
In contrast, our work targets novel view synthesis given only a single input image and a target camera pose. This setting spans both short and long-range viewpoint changes. Under such conditions, methods such as NeRF [30] and 3DGS [23] struggle to extrapolate effectively from a single image, even when augmented with generative guidance as in [42, 45].
Transformers: GeoGPT [36] was one of the early works to perform view synthesis using transformers. NViST [22] adopts a transformer-based encoder-decoder architecture [48, 9] to predict a radiance field from a single image, enabling novel view synthesis via NeRF-style volumetric rendering. However, NViST suffers from loss of fine details due to aggressive downsampling (by a factor of 12), and it struggles to synthesize long-range viewpoints (when the target frame is more than 15 frames away from the input frame in a 30‑frame sweep).
@@ 2 Related Work
Pretrained Diffusion Models: MVDiffusion [43], Zero123++ [39], SyncDreamer [27], Wonder3D [29], EpiDiff [21], BoostDream [62], MVDiff [3], CAT3D [14], Magic-Boost [57], Cycle3D [44], GenWarp [38], use a pretrained or finetuned diffusion model. MVDiffusion modifies the Stable Diffusion architecture by introducing a cross-branch attention mechanism, known as correspondence-aware attention (CAA), to model inter-view dependencies. SyncDreamer constructs a view frustum feature volume from all the target noisy views and injects these into the pretrained denoising Unet using depth-wise attention layers. EpiDiff [21] integrates an attention module guided by epipolar constraints into the intermediate and decoding stages of the U-Net, enabling the model to capture generalized epipolar geometry across views. GenWarp [38] introduces a warp and inpaint technique.
@@
Training Diffusion Model from Scratch: Several recent works train diffusion models from scratch for novel view synthesis, including Tseng et al. [46], Photometric-NVS [60], DiffDreamer [5], GIBR [1], and [17]. Photometric-NVS [60] introduces a two-stream latent diffusion architecture that independently processes the source and noisy target views, while exchanging information via pose-conditioned cross-attention mechanisms. GIBR [1] models 3D scenes using IB-planes and trains the diffusion process directly in pixel space, enabling learning of a joint distribution over multi-view observations and camera poses.
@@ 3 Method
3.1 Spectral Behavior of Diffusion
In [24, 11], authors study the spectral behavior of diffusion. The forward diffusion [19] process is given by:
@@ 3.1 Spectral Behavior of Diffusion
where $\mathbf{x}_{0}$ is the clean latent, $\mathbf{x}_{t}$ is the latent corresponding to the timestep $t$ and $\bar{\alpha}_{t}$ is the scaling factor. High-frequency components, representing fine details, are degraded more rapidly and prior to low-frequency components during the forward diffusion process [24, 11]. This property is also consistent in the reverse process of diffusion.
As shown in Choi et al. [7], diffusion models inherently favor lower frequencies, which implies that more emphasis must be placed on modelling high-frequency details. In addition, the noise component is often observed to deviate from a standard multivariate Gaussian distribution [41]. At later iterations of inversion, the noise encapsulates high-frequency information of the image and is high in variance. The signal variance decreases with inversion time, and the predicted noise’s variance increases with inversion time. The effective DDIM Inversion[40] iteration is:
$\mathbf{x}_{t+1}=$ $\underbrace{\left(\mathbf{x}_{t}-\sqrt{1-\overline{\alpha}_{t}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\sqrt{\frac{\overline{\alpha}_{t+1}}{\overline{\alpha}_{t}}}}_{\text{signal / mean},\,\mathbf{z}^{\text{inv}}_{\mu,t+1}}$ (2)
$\hskip 40.00006pt+\underbrace{\sqrt{1-\overline{\alpha}_{t+1}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}_{\text{noise / variance},\,\mathbf{z}^{\text{inv}}_{\sigma,t+1}},$
where $\mathbf{x}_{t+1}$ is the noisy latent at timestep $t+1$. $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the predicted noise at timestep $t$, estimated by the diffusion U-Net during the reverse process.
@@
The signal/mean in Equation (2) is the coarse-grained image representation on which we perform the view translation. In addition, the noise/variance in Equation (2) encodes image-specific features that are recovered during the denoising process [41]. For the task of novel-view synthesis, this noise/variance can be used to induce high-frequency details into the view-transformed latent, which can then be fed to DDIM sampling. Based on the aforementioned discussion, we formalize two things for the task of novel view synthesis:
@@
3.2 DDIM-inverted Latents
Let $\mathbf{z}^{\text{inv}}_{t}$ be the DDIM-inverted latent. If we use this latent at $t=T$, we may see that the DDIM sampled image deviates from the input image, especially when we do it in fewer DDIM steps [2, 63, 12]. Thus, we fix $t$ = 600 and get our DDIM inverted noisy initial latent in 30 DDIM steps. Further, the signal/mean term $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ of Equation (2) is the diffusion network’s estimate of the clean latent which we obtain at $t$ by denoising $\mathbf{z}^{\text{inv}}_{t}$ according to the diffusion score model $\boldsymbol{\epsilon}_{\theta}$. This signal/mean term is what we feed into TUNet for view translation. We visualise the reconstructed image corresponding to signal/mean term in Figure 3. We observe that the reconstructed image primarily comprises of the low‑frequency components of the input image. Therefore, in order to impose high-frequency information, we make use of the noise/variance term $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ from Equation (2) that re‑injects the predicted noise at the level $t$. We utilize a pretrained latent diffusion model (LDM) [35] as our generative prior, which we use to compute the DDIM-inverted latents. We first describe the TUNet model, followed by the Fusion Strategy. As we fix timestep $t$ at 600, we drop the subscript $t$ while representing signal/mean and noise/variance terms: $\mathbf{z}^{\text{inv}}_{\mu,t+1}$ and $\mathbf{z}^{\text{inv}}_{\sigma,t+1}$ in Equation (2) from now on.
Figure 3: Mean of the DDIM inverted latent at $t=400,600,800$, respectively. Latent is decoded using VAE for visualization. Original $\times512$ image. At $t=400$, the mean reflects dominant low frequencies which precludes generation of diverse images. At $t=800$, the low frequency component is extremely weak. $t=600$ provides a weak yet effective signal for translation.
3.3 TUNET Architecture
TUNet is a U-Net  [37] inspired encoder-decoder architecture designed to predict the DDIM-inverted latent’s mean representation of the target view. TUNet introduces cross attention between an input or reference view and a target view, enabling effective feature transfer between viewpoints. The architecture is conditioned on both camera parameters and class embeddings at multiple stages to preserve geometric consistency and semantic integrity.
3.3.1 Input and Conditioning
@@
3.3.2 Encoder (Down Blocks)
@@
3.3.3 Bottleneck and Decoder (Mid + Up Blocks)
@@
3.3.4 Cross-Attention Module
A cross-attention mechanism is integrated in the mid and up blocks, enabling information flow from the reference to the target view using ray information and latent feature alignment. Let $\mathbf{r_{\text{ref}}}$ denote the ray embeddings of the reference view and $\mathbf{r_{\text{tar}}}$ denote the ray embeddings of the target view. We use standard ray parameterization as in NeRF [30] to compute ray origins and directions for camera pose encoding to get $\mathbf{r_{\text{ref}}}$ and $\mathbf{r_{\text{tar}}}$. Let $\mathbf{z^{\text{inv}}_{\text{ref},\mu}}$ be the DDIM-inverted latent mean of the reference image, and $\mathbf{f_{\text{tar}}}$ be the intermediate target feature maps at the cross-attention block.